Troy Engines, Ltd., manufactures a variety of engines for use in heavy equipment. The company has always produced all of the necessary parts for its engines, including all of the carburetors. An outside supplier has offered to sell one type of carburetor to Troy Engines, Ltd., for a cost of $36 per unit. To evaluate this offer, Troy Engines, Ltd., has gathered the following information relating to its own cost of producing the carburetor internally:

Per Unit 20,000 Units
Per Year
Direct materials $ 17 $ 340,000
Direct labor 10 200,000
Variable manufacturing overhead 2 40,000
Fixed manufacturing overhead, traceable 9 * 180,000
Fixed manufacturing overhead, allocated 12 240,000
Total cost $ 50 $ 1,000,000

*One-third supervisory salaries; two-thirds depreciation of special equipment (no resale value).

Required:

1. Assuming the company has no alternative use for the facilities that are now being used to produce the carburetors, what would be the financial advantage (disadvantage) of buying 20,000 carburetors from the outside supplier?

2. Should the outside supplier’s offer be accepted?

3. Suppose that if the carburetors were purchased, Troy Engines, Ltd., could use the freed capacity to launch a new product. The segment margin of the new product would be $200,000 per year. Given this new assumption, what would be the financial advantage (disadvantage) of buying 20,000 carburetors from the outside supplier?

4. Given the new assumption in requirement 3, should the outside supplier’s offer be accepted?
Answer #1
1
Per unit Total 20000 units
Make Buy Make Buy
Direct materials 17 340000
Direct labor 10 200000
Variable manufacturing overhead 2 40000
Fixed manufacturing overhead traceable 3 60000
Purchase cost 36 720000
Total 640000 720000
Difference in favor of making = 640000-720000 = $80000
Financial (disadvantage) $(80000)
2
NO, Reject the offer
3
Make Buy
Total cost 640000 720000
Opportunity cost 200000
Total relevant cost 840000 720000
Difference in favor of buying = 840000-720000 = $120000
Financial advantage $120000
4
Yes, Accept the offer